OpenStudy (anonymous):

(-2y/3) -(23/6)=3/2.

OpenStudy (anonymous):

find the value of y with steps

OpenStudy (anonymous):

awww, ur so little

OpenStudy (anonymous):

go to www.wolframalpha.com and enter it like the other one m8.

OpenStudy (anonymous):

they are not giving the explantaion

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

google

jhonyy9 (jhonyy9):

-4y-23=9 -32=4y y=-32/4 y=-8